OpenStudy (anonymous):

steps to solving (log base 2 of 4)+(log base 3 of (1/9)) please!

OpenStudy (anonymous):

2^2=4 n 3^3=9. so for first one u get 2

OpenStudy (anonymous):

for first part it will be 2(exponent of 2 which gives 4) and for second part, the exponent of 3 that gives 1/9 is -2 so 2-2=0 @jennyneedhelp

OpenStudy (anonymous):

\[\log_{2} 4+\log_{3} 1/9\]can be written as\[\log_{2} 2^2+\log_{3} 3^-2\] now apply formula \[\log_{q} q^p=p \log_{q} q=p\]by using above formula we get \[2-2=0\]